# MMedPO: Aligning Medical Vision-Language Models with Clinical-Aware Multimodal Preference Optimization

### v1

@@ 4.2 Main Results
Comparison with Baseline Methods Enhanced by SFT. To demonstrate the compatibility of our approach with other training methods, we conduct experiments by applying MMedPO and other baseline methods to supervised fine-tuning (SFT). As shown in Table 1, MMedPO consistently outperforms the SFT baseline across all four datasets on both the Med-VQA and report generation tasks, with an average improvement of 14.2%. When compared to other baselines applied to SFT, MMedPO achieves significantly better performance, with an average improvement of 10.5% across all datasets. These results further corroborate the effectiveness and compatibility of our approach, demonstrating its ability to integrate seamlessly with other training techniques to enhance model alignment.
@@ 4.3 Quantitative Analysis
In this section, we first conduct ablation study to analyze the effectiveness of each strategy and component in MMedPO for enhancing factual accuracy. Then we evaluate the model’s compatibility based on different backbones and RAG technique. We further explore how our approach improves Med-LVLMs’ responses in terms of clinical significance and visual understanding.
@@ 4.3.1 Ablation Study
Ablation Study on Different Preference Curation Strategies. To assess the impact of different preference curation strategies in MMedPO, namely generating hallucinated medical responses and adding noise to localized lesion regions, we evaluated their performance on these two components. The results, presented in Figure 3, reveal that adding noise to localized lesion regions has a more pronounced effect on open-ended generation tasks (e.g., report generation) compared to generating hallucinated medical responses. For medical VQA tasks, the performance improvements from both preference curation processes are comparable. By integrating both strategies, MMedPO achieves the best overall performance across four datasets, effectively combining their strengths to maximize performance gains.
@@
Ablation Study on Clinical Relevance Score. To investigate the role of clinical relevance score as weight in the preference optimization process, we compare the results of applying this weight versus not applying it under different preference curation strategies. The results indicate that incorporating clinical relevance scores as weights in preference optimization improves the effectiveness of preference optimization. Specifically, as shown in Table 2, on the Med-VQA datasets, models utilizing clinical relevance scores as weights consistently outperform those without them, with an average improvement of 2.3%. Also significant performance gains are observed on the report generation task, where clinical relevance scores contributed positively across different preference curation strategies, achieving a clear average margin of 18.5%. The clinical relevance scores assigned to each preference pair provide positive benefits to preference optimization, helping the Med-LVLMs generate responses that are both more clinically meaningful and accurate.
Table 2: Comparison of performance across different datasets with and without clinical relevance score (CRS) for different preference curation strategies during preference optimization. Here, stage 1 and stage 2 denotes generating hallucinated medical responses and adding noise to localized lesion regions, respectively. We report the average score on each dataset.
@@
4.3.2 Does Collaboration Among Multiple Med-LLMs Outperform Using a Single Med-LLM?
To explore the impact of multi-agent collaboration mechanism in generating clinical relevance scores, we conduct analytical experiments on four datasets, comparing the performance using clinical relevance scores from single Med-LLM and multiple Med-LLMs. As shown in Table 3, we find that the consensus scores reached by multiple Med-LLMs positively contributes to performance improvement by an average of 3.6% over all datasets. This aligns with our expectations, as relying on a single Med-LLM can introduce biases. The observed improvement is driven by reduced bias through the collaborative efforts of multiple Med-LLMs, resulting in more accurate and clinically meaningful relevance evaluations. In addition, the performance gains on the Med-VQA task using multiple Med-LLMs are notably larger compared to the report generation task. This may be attributed to greater disagreement among Med-LLMs on rejected VQA answers, allowing them to benefit more from achieving consensus.
Table 3: Comparison of model performance using clinical relevance scores from single Med-LLM and multiple Med-LLMs for MMedPOẆe report the average score on each dataset.
@@ 4.3.2 Does Collaboration Among Multiple Med-LLMs Outperform Using a Single Med-LLM?
4.3.3 Impact of Localized Lesion Noise in MMedPO
Table 4: Performance comparison between introducing local noise and global noise on the stage of constructing preference data by adding noise to medical images. We report the average score on each dataset.
@@ 4.3.4 Compatibility Analysis
To evaluate the compatibility of our approach with different base models, particularly more powerful backbone architectures, we replace the backbone of LLaVA-Med-1.5 and conduct a series of experiments based on this configuration. Specifically, we apply our method to LLaVA-Med++ [54], which uses LLaMA-3 [11] as language backbone and enhances its performance using a large-scale medical multimodal dataset MedTrinity-25M. As illustrated in Table 4, similar to the results obtain with LLaVA-Med-1.5, applying MMedPO leads to performance improvements across all four datasets. These findings highlight the strong compatibility and effectiveness of our approach when integrated with other powerful Med-LVLMs. MMedPO can be transferred to a wider range of base models, demonstrating strong generalizability for applications in clinical scenarios.
@@ 4.4 Qualitative Analysis and Case Study
In this section, we further conduct qualitative experiments and case analyses to demonstrate the effectiveness of MMedPO.
@@ 4.4.1 Qualitative Analysis
How does MMedPO in Improving Visual Understanding? To better understand the model’s visual comprehension capability, we visualize its attention map on image tokens. As shown in Figure 5, compared to the attention map of the original LLaVA-Med-1.5, the utilization of MMedPO significantly enhances the model’s focus on visual information, particularly on critical lesion areas. This allows the model to extract sufficient information from visual inputs and improve consistency between text and images. Thus the model can reduce hallucinations and provide more accurate answers.
Figure 5: Visualization of attention map of image tokens. The red box region is labeled with the attentions that can be enhanced by MMedPO.
Analysis Clinical Significance of Model’s Response. Through the analysis of previous results, Med-LVLMs enhanced by MMedPO demonstrate a significant improvement in the accuracy of generated responses. Additionally, from the clinical perspective, we aim to evaluate the clinical significance of the responses to verify the effectiveness of MMedPO in enhancing the clinical relevance of the model’s outputs. As demonstrated in Figure 6, on this case, the Med-LVLMs with MMedPO outperforms both the original Med-LVLM and the one applied with DPO method. The response with MMedPO accurately capture the condition of the cardiac silhouette and rib fracture in the image, aligning with the ground truth. Additionally, the response with MMedPOimproves clinical significance judged by Med-LLMs. The original model and other baselines, in contrast, produce repetitive and clinically irrelevant contents, leading to inferior responses. The evaluation of response using clinical relevance from Med-LLMs quantitatively shows that MMedPO consistently achieves significantly higher clinical relevance scores.
Figure 6: Examples demonstrating the clinical relevance of responses generated by MMedPO. Our approach not only enhances the factual accuracy of the model’s generated responses but also significantly improves their clinical relevance, including various meaningful medical-level explanations.
@@ 4.4.2 Case Study
We analyze two examples from Medical VQA task to illustrate how the model fine-tuned with MMedPO reduces factuality errors. As illustrated in Figure 7, the model fine-tuned using the MMedPO method shows improved performance in factual accuracy. For example, when asked about pathology, the model provides a more detailed response, focusing on the problem of renal cyst, which is similar to the ground truth, outperforming both LLaVA-Med and LLaVA-Med with DPO fine-tuning. In another case, when the user inquires about the patient’s condition, MMedPO accurately identifies hydrocephalus based on the image, while the other two models fail to do so. This demonstrates that MMedPO effectively reduces hallucinations in Med-LVLMs, minimizing factual errors in multimodal understanding tasks.
Figure 7: Illustrations of factuality enhancement by MMedPO.
@@ 5 Related Work
Factuality Issues in Med-LVLMs. The development of Large Vision-Language Models (LVLMs) is progressing rapidly [28, 29, 65, 55, 56, 50, 21], which has, in turn, driven advancements in Medical Vision-Language Models (Med-LVLMs), achieving promising results in the medical field [22, 30, 37, 48, 53, 52]. However, the current Med-LVLMs still exhibit significant factual errors [46, 24, 49, 34, 5, 16]. For example, they often lack sufficient judgment ability for complex content, and frequently generate responses with hallucinations that contradicts the visual information provided. This issue is particularly pronounced in medical domain, as it can potentially lead to misdiagnoses or missed diagnoses. Recently, there are several benchmarks [49, 34] that highlight the factuality issues of Med-LVLMs on multiple tasks such as the visual question answering and report generation.
Preference Optimization in Med-LVLMs. Aligning with human preferences for large models is an effective way to address hallucination issues [63, 20, 45, 64, 10, 8, 38]. Preference fine-tuning in LVLMs generally involves two approaches: one aligns models based on human feedback [1, 33], while the other uses feedback generated by AI [20, 63, 64, 59, 45, 44, 8, 61]. Recently, the preference fine-tuning technique has also been adapted for medical imaging [2, 35, 12] by generating dispreferred responses using GPT-4 or the target Med-LVLM. Although these methods have shown promise, they neglect the clinical relevance of the generated samples. Moreover, in Med-LVLMs, visual information from local regions plays a pivotal role in generating accurate responses. However, these methods rarely guide the model’s focus toward specific lesion regions during preference fine-tuning. To address these challenges, our work incorporates quantified clinical relevance scores as weights to enhance preference fine-tuning. Additionally, we introduce localized noise into medical images to construct dispreference, which directs the model’s attention to critical local regions, thereby improving its visual understanding of important lesions.
@@ 6 Conclusion
In this work, we propose a clinical-aware multimodal preference optimization approach, which considers the clinical relevance of each preference sample in preference optimization. This method enhances alignment of Med-LVLMs while effectively reducing factual errors. Specifically, to construct multimodal preference data, we introduce plausible hallucinations through target Med-LVLMs or GPT-4o and apply local noise to critical lesion regions. Furthermore, we assign clinical relevance for data samples through Med-LLMs and visual tools, and then incorporate these scores as weights in the preference fine-tuning process. We evaluate the effectiveness of MMedPO on the Med-VQA and report generation tasks, demonstrating superior performance.
@@ A.1 Data Statistics
Table 5: Data statistics for the training set of four datasets under two different task settings. “Train (stage 2)” refers to the number of preference data introducing noise to localized lesion regions, while “Train (stage 1)” indicates the number of preference data with hallucinated medical responses.
Dataset Train (stage 2) Train (stage 1) Train (all)
@@ A.2 Involved Datasets
We leverage four open-source medical vision-language datasets: MIMIC-CXR [17], IU-Xray [9], SLAKE [26], and VQA-RAD [19]. These datasets are designed for different tasks: the first two focus on medical report generation, while the latter two are tailored for medical visual question answering.
@@ Appendix B Hyperparameter Settings
For the usage of visual tools, we employ “disease” as the text description to guide MedKLIP [47] in generating heatmaps. For multi-agent collaboration, the process is conducted over 5 rounds. During score normalization, the parameters are set as: $\alpha=0.75$, $\beta=1.25$, $\mu=1$, and $\sigma^{2}=0.1$. All experiments are implemented using PyTorch 2.1.2 on four NVIDIA RTX A6000 GPUs, with training requiring approximately 2 to 3 hours.
@@ Appendix C Involved Baselines
  • DPO [33] is a fine-tuning approach designed to align large language models (LLMs) with human preferences in a stable, efficient, and computationally lightweight manner. Unlike traditional Reinforcement Learning from Human Feedback (RLHF), which involves training a reward model and using reinforcement learning to maximize the reward, DPO simplifies the process by reframing the problem. It parameterizes the reward model in a way that allows the optimal policy to be derived directly through a classification loss, eliminating the need for complex sampling or extensive hyperparameter tuning during fine-tuning.
  • Self-Rewarding [58] is a novel approach where the language model itself acts as a judge, generating rewards via LLM-as-a-Judge prompting during training. Unlike traditional methods that rely on reward models trained from human preferences, which are limited by human performance and static design, this method enables the model to iteratively improve both its instruction-following abilities and its reward-generating quality during iterative DPO training.
  • STLLaVA-Med [35] refines the preference selection process using GPT-4o and applies it in medical vision-language tasks. STLLaVA-Med extends the DPO approach by incorporating a self-training mechanism specifically tailored for the medical domain.

### v2

@@ 4.2 Main Results
Comparison with Baseline Methods Enhanced by SFT. To demonstrate the compatibility of our approach with other training methods, we conduct experiments by applying MMedPO and other baseline methods to SFT. As shown in Table 1, MMedPO consistently outperforms the SFT baseline across all four datasets, with an average improvement of 14.2%. When compared to other baselines applied to SFT, MMedPO achieves significantly better performance, with an average improvement of 10.5%. These results further corroborate the effectiveness and compatibility of our approach, demonstrating its ability to integrate seamlessly with other training techniques to enhance model alignment.
@@ 4.3 Quantitative Analysis
In this section, we first conduct ablation study to analyze the effectiveness of each strategy and component in MMedPO for enhancing factual accuracy. Then, we evaluate the model’s compatibility with different backbones. We further explore how our approach improves Med-LVLMs’ responses in terms of clinical significance and visual understanding.
@@ 4.3.1 Ablation Study
Different Preference Curation Strategies. To assess the impact of different preference curation strategies in MMedPO, namely generating hallucinated medical responses and adding noise to localized lesion regions, we evaluated their performance on these two components. The results, presented in Figure 3, reveal that adding noise to localized lesion regions has a more pronounced effect on open-ended generation tasks (e.g., report generation) compared to generating hallucinated medical responses. For medical VQA tasks, the performance improvements from both preference curation processes are comparable. By integrating both strategies, MMedPO achieves the best overall performance across four datasets, effectively combining their strengths to maximize performance gains.
@@
Table 2: Comparison of performance across different datasets with and without clinical relevance score (CRS) for different preference curation strategies. Here, stage 1 and stage 2 denote generating hallucinated medical responses and adding noise to localized lesion regions, respectively. We report the average score on each dataset.
@@
Clinical Relevance Score. To investigate the role of clinical relevance score as weight in the preference optimization process, we compare the results of applying this weight versus not applying it under different preference curation strategies. The results indicate that incorporating clinical relevance scores as weights in preference optimization improves the effectiveness of fine-tuning. Specifically, as shown in Table 2, for VQA task, models utilizing clinical relevance scores as weights consistently outperform those without them, with an average improvement of 2.3%. Also, significant performance gains are observed on the report generation task, where clinical relevance scores contributed positively across different preference curation strategies, achieving a clear average margin of 18.5%. The clinical relevance scores assigned to each preference pair provide positive benefits to preference optimization, helping the Med-LVLMs generate responses that are more clinically meaningful and accurate.
4.3.2 Multiple vs. Single Med-LLM
To explore the impact of the multi-agent collaboration mechanism in generating clinical relevance scores, we conduct analytical experiments, comparing the performance using clinical relevance scores from single Med-LLM and multiple Med-LLMs. As shown in Table 3, we find that the consensus scores reached by multiple Med-LLMs positively contribute to performance improvement by an average of 3.6% over four datasets. This aligns with our expectations, as relying on a single Med-LLM will introduce biases. The observed improvement is driven by reduced bias through the collaborative efforts of multiple Med-LLMs, resulting in more accurate and clinically meaningful relevance evaluations. In addition, the performance gains on the Med-VQA task using multiple Med-LLMs are notably larger compared to the report generation task. This may be attributed to greater disagreement among Med-LLMs on rejected VQA answers, allowing them to benefit more from achieving consensus.
4.3.3 Impact of Localized Lesion Noise
Table 3: Comparison of model performance using clinical relevance scores from single Med-LLM and multiple Med-LLMs for MMedPO. We report the average score on each dataset.
@@ 4.3.3 Impact of Localized Lesion Noise
Table 4: Performance comparison between introducing local noise and global noise on the stage of constructing preference data by adding noise to medical images.
@@ 4.3.4 Compatibility Analysis
To evaluate the compatibility of our approach with different base models, particularly more powerful backbone architectures, we replace the backbone of LLaVA-Med-1.5 and conduct a series of experiments based on this configuration. Specifically, we apply our method to LLaVA-Med++ (Xie et al., 2024), which uses LLaMA-3 (Dubey et al., 2024) as language backbone and enhances its performance using a large-scale medical multimodal dataset MedTrinity-25M. As illustrated in Table 4, similar to the results obtain with LLaVA-Med-1.5, applying MMedPO leads to performance improvements across all four datasets. These findings highlight the strong compatibility and effectiveness of our approach when integrated with other powerful Med-LVLMs. MMedPO can be transferred to a wider range of base models, demonstrating strong generalizability for applications in clinical scenarios.
@@ 4.4 Qualitative Analysis and Case Study
In this section, we further conduct qualitative experiments and case analyses.
@@ 4.4.1 Qualitative Analysis
How does MMedPO in Improving Visual Understanding? To better understand the model’s visual comprehension capability, we visualize its attention map on image tokens. As shown in Figure 5, compared to the attention map of the original model, the utilization of MMedPO significantly enhances the model’s focus on visual information, particularly on critical lesion areas. This allows the model to extract sufficient information from visual inputs and improve consistency between text and images. Thus the model can reduce hallucinations and provide more accurate answers.
Figure 5: Visualization of attention map of image tokens. The red box region is labeled with the attentions that are enhanced.
Figure 6: Examples demonstrating the clinical relevance of responses generated by MMedPO. Our approach not only enhances the factual accuracy but also significantly improves the clinical relevance, including various meaningful medical-level explanations.
Analysis Clinical Significance of Model’s Response. Through the analysis of previous results, Med-LVLMs enhanced by MMedPO demonstrate a significant improvement in factuality accuracy. Additionally, from the clinical perspective, we aim to evaluate the clinical significance of the responses to verify the effectiveness of MMedPO in enhancing the clinical relevance of the model’s outputs. As demonstrated in Figure 6, Med-LVLMs with MMedPO outperforms both the original model and the one applied with DPO. The response with MMedPO accurately capture the condition of the cardiac silhouette and rib fracture in the image, aligning with the ground truth. This also improves clinical significance judged by Med-LLMs, whereas the original model and other baselines produced duplicate and clinically irrelevant content. The evaluation of response using clinical relevance from Med-LLMs quantitatively shows that MMedPO consistently achieves significantly higher clinical relevance scores.
Figure 7: Illustration of factuality enhancement by MMedPO.
@@ 4.4.2 Case Study
We analyze two examples from Medical VQA task to illustrate how the model fine-tuned with MMedPO reduces factuality errors. As illustrated in Figure 7, MMedPO shows improved performance in factual accuracy. In this case, when asked about pathology, MMedPO provides a more detailed response, focusing on the problem of cyst, which is similar to the ground truth, outperforming both LLaVA-Med and LLaVA-Med with DPO. This demonstrates that MMedPO effectively reduces hallucinations in Med-LVLMs, minimizing factual errors in multimodal understanding tasks.
@@ 5 Related Work
Factuality Issues in Med-LVLMs. The development of Large Vision-Language Models (LVLMs) is progressing rapidly (Liu et al., 2024a, b; Zhu et al., 2023; Bai et al., 2023; Xia et al., 2025, 2024d; Han et al., 2025), which has, in turn, driven advancements in Medical Vision-Language Models (Med-LVLMs), achieving promising results in the medical field (Li et al., 2023a; Moor et al., 2023; Thawkar et al., 2023; Wu et al., 2023c). However, the current Med-LVLMs still exhibit significant factual errors (Wu et al., 2023a; Li et al., 2023b; Xia et al., 2024a; Chen et al., 2024; Jiang et al., 2024). For example, they often lack sufficient judgment ability for complex content, and frequently generate responses with hallucinations that contradicts the visual information provided. This issue is particularly pronounced in medical domain, as it can potentially lead to misdiagnoses or missed diagnoses. Recently, there are several benchmarks (Xia et al., 2024a; Royer et al., 2024) that highlight the factuality issues of Med-LVLMs on multiple tasks such as the visual question answering and report generation.
Preference Optimization in Med-LVLMs. Aligning with human preferences for large models is an effective way to address hallucination issues (Lee et al., 2024; Zhou et al., 2024a, b; Deng et al., 2024). Preference fine-tuning in LVLMs generally involves two approaches: one aligns models based on human feedback (Bai et al., 2022; Rafailov et al., 2023), while the other uses feedback generated by AI (Lee et al., 2024; Zhou et al., 2024a, b; Wang et al., 2024a; Zhou et al., 2025; Tong et al., 2025). Recently, the preference fine-tuning technique has also been adapted for medical imaging (Banerjee et al., 2024; Sun et al., 2024; Hein et al., 2024) by generating dispreferred responses using GPT-4 or the target Med-LVLM. Although these methods have shown promise, they neglect the clinical relevance of generated samples. In Med-LVLMs, local visual information is crucial for accurate responses, yet current approaches rarely guide the model’s focus to specific lesion areas during preference fine-tuning. To tackle these issues, we incorporate quantified clinical relevance scores as weights to enhance modality alignment and introduce localized noise in medical images to construct dispreference, improving its understanding of key lesions.
@@ 6 Conclusion
In this work, we propose a novel clinical-aware multimodal preference optimization approach named MMedPO  which considers the clinical relevance of each preference sample in preference optimization process. This method enhances Med-LVLM alignment while effectively reducing factual hallucinations. Specifically, to construct multimodal preference data, we introduce plausible hallucinations and apply local noise to critical lesion regions. Furthermore, we assign clinical relevance for data samples through Med-LLMs and visual tools, and then incorporate these scores as weights in the preference fine-tuning process. We evaluate the effectiveness of MMedPO on the Med-VQA and report generation tasks, demonstrating superior performance.
@@ Acknowledgement
Impact Statement
The broader impact of this work lies in its potential to enhance the reliability and accuracy of AI-driven medical diagnostics by reducing hallucinations and improving visual-textual alignment in Med-LVLMs. This advancement could lead to more trustworthy AI tools in healthcare, benefiting patient outcomes. However, ethical considerations are crucial to ensure responsible deployment, prevent misuse, and avoid over-reliance on AI-generated medical advice. Future societal benefits may include reduced diagnostic errors and improved healthcare efficiency, but ongoing research and ethical oversight are essential to align these advancements with the best interests of patients and providers.
@@ A.1 Data Statistics
Table 5: Data statistics for the training set of four datasets under two different task settings. “Train (visual)” refers to the number of visual-only preference data, while “Train (text)” indicates the number of text-only preference data.
Dataset Train (visual) Train (text) Train (all)
@@ A.2 Involved Datasets
We leverage four open-source medical vision-language datasets: MIMIC-CXR (Johnson et al., 2020), IU-Xray (Demner-Fushman et al., 2016), SLAKE (Liu et al., 2021), and VQA-RAD (Lau et al., 2018). These datasets are designed for different tasks: the first two focus on medical report generation, while the latter two are tailored for medical visual question answering.
@@ Appendix B Hyperparameter Settings
For the usage of visual tools, we employ “disease” as the text description to guide MedKLIP (Wu et al., 2023b) in generating heatmaps. For multi-agent collaboration, the process is conducted over 5 rounds . During score normalization, the parameters are set as: $\alpha=0.75$, $\beta=1.25$, $\mu=1$, and $\sigma^{2}=0.1$. All hyperparameters are kept consistent across the experiments to eliminate any potential bias introduced by hyperparameter tuning. All experiments are implemented using PyTorch 2.1.2 on four NVIDIA RTX A6000 GPUs, with training requiring approximately 2 to 3 hours.
Models IU-Xray MIMIC-CXR
BL-1 BL-2 BL-3 BL-4 ROUGE-L METEOR BL-1 BL-2 BL-3 BL-4 ROUGE-L METEOR
LLaVA-Med v1.5 38.42 13.40 4.74 1.67 10.31 10.95 29.41 10.19 3.58 1.26 9.38 7.71
+ Self-Rewarding 38.25 13.17 3.61 1.08 10.38 10.52 29.29 10.32 3.67 1.30 9.27 7.73
+ DPO 41.63 15.13 5.56 2.03 12.95 17.13 29.61 10.29 3.61 1.27 9.45 7.81
+ POVID 50.84 20.65 8.38 3.31 24.33 30.05 29.68 10.29 3.61 1.26 9.66 7.84
+ SIMA 42.67 16.82 5.98 2.95 22.87 29.10 29.58 10.23 3.59 1.24 9.58 7.49
+ FiSAO 51.10 20.92 8.64 3.59 25.72 30.82 29.76 10.37 3.74 1.39 9.68 7.62
+ STLLaVA-Med 42.38 15.27 5.59 1.20 10.58 10.51 29.33 10.27 3.58 1.27 9.29 7.72
+ MMedPO (Ours) 55.36 23.85 10.34 4.39 29.52 34.16 33.67 11.91 4.28 1.54 11.13 10.03
Models IU-Xray MIMIC-CXR
BL-1 BL-2 BL-3 BL-4 ROUGE-L METEOR BL-1 BL-2 BL-3 BL-4 ROUGE-L METEOR
+ Stage 1 (Single-LLM) 43.45 16.05 5.99 2.21 19.66 22.65 29.41 10.19 3.58 1.26 9.33 7.77
+ Stage 1 (Multi-LLMs) 43.95 16.44 6.21 2.31 19.57 22.92 29.85 10.38 3.65 1.28 9.62 8.18
+ Stage 2 30.15 10.86 4.53 1.40 29.02 34.26 30.96 10.89 3.87 1.38 9.85 8.81
+ Stage 1+2 (Single-LLM) 30.29 11.16 4.78 1.55 29.30 34.22 32.96 11.63 4.14 1.46 10.99 10.03
+ Stage 1+2 (Multi-LLMs) 30.33 11.23 4.76 1.54 29.52 34.16 33.67 11.91 4.28 1.54 11.13 10.05
@@ Appendix C Involved Baselines
  • DPO (Rafailov et al., 2023) is a fine-tuning approach designed to align large language models (LLMs) with human preferences in a stable, efficient, and computationally lightweight manner. Unlike traditional Reinforcement Learning from Human Feedback (RLHF), which involves training a reward model and using reinforcement learning to maximize the reward, DPO simplifies the process by reframing the problem. It parameterizes the reward model in a way that allows the optimal policy to be derived directly through a classification loss, eliminating the need for complex sampling or extensive hyperparameter tuning during fine-tuning.
  • Self-Rewarding (Yuan et al., 2024) is a novel approach where the language model itself acts as a judge, generating rewards via LLM-as-a-Judge prompting during training. Unlike traditional methods that rely on reward models trained from human preferences, which are limited by human performance and static design, this method enables the model to iteratively improve both its instruction-following abilities and its reward-generating quality during iterative DPO training.
  • STLLaVA-Med (Sun et al., 2024) refines the preference selection process using GPT-4o and applies it in medical vision-language tasks. STLLaVA-Med extends the DPO approach by incorporating a self-training mechanism specifically tailored for the medical domain.
  • POVID (Zhou et al., 2024a) addresses the hallucination problem in vision-language models by generating feedback data using AI models. It uses ground-truth instructions as preferred responses and creates dispreferred data by injecting plausible hallucinations and distorting images, integrating these strategies into an RLHF pipeline via DPO.
  • FiSAO (Cui et al., 2024) introduces a fine-grained self-alignment optimization method that leverages the model’s own visual encoder to improve vision-language alignment. By utilizing token-level feedback from the vision encoder, it enhances alignment without the need for additional external data, outperforming traditional preference tuning methods.
  • SIMA (Wang et al., 2024b) is a framework that enhances visual and language modality alignment through self-improvement, eliminating the need for external models or data. It uses prompts from existing datasets to self-generate responses and employs an in-context self-critic mechanism with vision metrics to select optimal response pairs for preference tuning.